# Visual Autoregressive Modeling for Instruction-Guided Image Editing

### v1

@@ Introduction
The critical challenge of scale mismatch introduced by the efficient finest-scale approach raises a fundamental question: which source scales are truly necessary for high-fidelity editing? To investigate such scale dependencies of target residuals and source residuals, we perform a diagnostic analysis of the self-attention mechanisms in a model trained on the full-scale source features. This full-scale setting allows the model to freely attend to all source scales.
Our empirical investigation of a representative example in Figure 3 reveals distinct attention patterns for different Transformer layers. In the first self-attention layer, when predicting tokens for a given target scale, the attention mechanism distributes broadly, focusing heavily on the corresponding and all coarser source scales. This pattern indicates that the initial layer is responsible for establishing the global layout and long-range dependencies. This behavior shifts in deeper layers, where attention patterns become highly localized. They exhibit strong diagonal structures, suggesting that attention is primarily confined to tokens in the spatial neighborhood. This functional transition suggests a shift from global structuring to local refinement, where the model copies and adjusts high-frequency details. For this latter task, the fine-grained information provided by $\mathbf{F}_{K}^{(src)}$ is sufficient. This motivates us to design a hybrid solution that provides the scale-aligned reference in the first layer and all subsequent layers only attend to the finest-scale source.
@@
Figure 3: Self-attention score heatmaps of the full-scale-based transformer among different layers in one edit. The dependency patterns are different in the first self-attention layer and others, inspiring a scale-aligned reference module as an improvement.
Scale-Aligned Reference
@@
During generation, when predicting the tokens for target scale $k$, the model’s first self-attention layer computes queries $\mathbf{Q}_{k}^{(tgt)}$ and attends to a combined set of keys and values, which are derived from two sources: (i) the newly generated, scale-aligned reference feature $\mathbf{F}_{k}^{(ref)}$ and (ii) the causal history of previously generated target tokens. Specifically, the attention output $\hat{\mathbf{O}}_{k}^{(tgt)}$ is computed as:
@@
$\hat{\mathbf{O}}_{k}^{(tgt)}=$ $\operatorname{Softmax}\left(\frac{\mathbf{Q}_{k}^{(tgt)\top}\left[\mathbf{K}_{k}^{(ref)},\mathbf{K}_{1}^{(tgt)},\cdots,\mathbf{K}_{k}^{(tgt)}\right]}{\sqrt{d}}\right)$ (7)
$\cdot$ $\left[\mathbf{V}_{k}^{(ref)\top},\mathbf{V}_{1}^{(tgt)\top},\cdots,\mathbf{V}_{k}^{(tgt)\top}\right]^{\top},$
where $(\mathbf{K}_{k}^{(ref)},\mathbf{V}_{k}^{(ref)})$ are projected from the scale-aligned source reference $\mathbf{F}_{k}^{(ref)}$, and $(\mathbf{K}_{1:k-1}^{(tgt)},\mathbf{V}_{1:k-1}^{(tgt)})$ are from all previously generated target tokens.
@@
Crucially, this SAR mechanism is applied only to the first self-attention layer. All subsequent layers operate using only finest-scale conditioning, attending to $\mathbf{F}_{K}^{(src)}$ and the target’s history residuals. This allows VAREdit to capture multi-scale dependencies while achieve the efficiency of the finest-scale approach.
Experiments
Experimental Setup
Datasets
@@
VAREdit is trained on a large-scale dataset of 3.92 million paired examples, which are aggregated from the SEED-Data-Edit (Ge et al. 2024) and ImgEdit (Ye et al. 2025) datasets. From the SEED-Data-Edit dataset, we first extracted all single-turn samples and proceeded to decompose multi-turn conversations into single-turn editing pairs. These generated pairs were then filtered using a vision-language model (Team et al. 2025) to remove instances with poor instruction-following quality. To finalize the training data, all single-turn samples from ImgEdit were also incorporated. Additional details regarding this data processing pipeline are available in the Appendix.
For evaluation, we comprehensively assess VAREdit on two established benchmarks: the EMU-Edit (Sheynin et al. 2024), which contains 3,589 examples across 8 distinct editing types, and the PIE-Bench (Ju et al. 2023), featuring 700 examples that cover 10 different editing types.
Evaluation Metrics
@@
Standard benchmarks like EMU-Edit and PIE-Bench rely on CLIP-based scores (Radford et al. 2021). EMU-Edit employs caption-image similarity (CLIP-Out.) and text-image directional similarity (CLIP-Dir.), whereas PIE-Bench measures similarity for the whole image (CLIP-Whole) and the edited region (CLIP-Edit). However, these metrics often fail to capture important editing quality aspects, such as spurious modifications or incomplete edits. To address these shortcomings, we also adopt the evaluation protocol from OmniEdit (Wei et al. 2025), which uses GPT-4o (Hurst et al. 2024) as an automated judge to provide two key scores on a scale of 0-10:
  • GPT-Success (Suc.): Measures the adherence to the editing instruction where higher is better.
  • GPT-Overedit (Over.): Assesses the preservation of unedited regions where higher is better.
Since a model could achieve a perfect GPT-Over. score by simply disregarding the edit instruction and outputting the original image, we introduce GPT-Balance (Bal.), the harmonic mean of GPT-Suc. and GPT-Over. metrics. This balanced score serves as our primary metric for overall editing performance. Detailed prompts and computation methods for these metrics are provided in the Appendix.
Compared Approaches
@@
To ensure a comprehensive and rigorous evaluation, we assess VAREdit against several state-of-the-art tuning-based methods. Our comparative analysis includes a broad spectrum of leading diffusion-based methods: InstructPix2Pix (Brooks et al. 2023), UltraEdit (Zhao et al. 2024a), OmniGen (Xiao et al. 2025), AnySD (Yu et al. 2025), ACE++ (Mao et al. 2025), and ICEdit (Zhang et al. 2025). Furthermore, we include EditAR (Mu et al. 2025), a vanilla AR method for benchmarking the effectiveness of our VAR-based framework. More details of these baselines are provided in the Appendix.
Implementation Details
@@
Our VAREdit models are initialized with weights from the pre-trained Infinity model (Han et al. 2025). To differentiate source image tokens from target image tokens, we introduce a positional offset of $\Delta=(64,64)$ to the 2D Rotary Position Embeddings (2D-RoPE) for all source tokens. To investigate scaling properties, we develop two distinct model sizes: VAREdit-2.2B and VAREdit-8.4B. The 2.2B model undergoes a two-stage training procedure: an initial 8k iterations trained at $256\times 256$ resolution with batch size 1,536 and learning rate 6e56e-5, followed by 7k fine-tuning iterations at $512\times 512$ with batch size 960 and learning rate 1.875e51.875e-5. The larger 8.4B model is trained directly at $512\times 512$ resolution for 26k iterations with batch size 1,536 and learning rate 6e56e-5. During inference, we employ a classifier-free guidance (CFG) strength of $\eta=4$ and a logits temperature of $\tau=0.5$. Please refer to the Appendix for more details.
@@
VAREdit 8.4B 0.276 0.115 7.512 7.519 6.773 0.255 0.224 8.156 7.490 7.298 1.2s
Table 1: Quantitative results of all methods on EMU-Edit and PIE-Bench benchmarks. The best results are marked in bold.
Figure 4: Fine-grained categorical GPT-Suc. and GPT-Bal. scores of all methods on EMU-Edit and PIE-Bench benchmarks.
Quantitative Results
The quantitative results in Table 1 demonstrate the superiority of VAREdit in both editing performance and efficiency.
Editing Quality
VAREdit consistently outperforms all diffusion-based and autoregressive baselines on our primary metric, GPT-Balance. Our 8.4B model achieves a GPT-Bal. score of 6.77 on EMU-Edit and 7.30 on PIE-Bench, surpassing the strongest competitor (ICEdit on EMU, UltraEdit on PIE) by 41.5% and 30.8%, respectively. This underscores VAREdit’s ability to perform precise edits while preserving unchanged regions. Notably, some methods like OmniGen achieve higher GPT-Over. scores by being overly conservative and failing to execute the requested edit, resulting in low GPT-Suc. and consequently poor GPT-Bal. scores. In contrast, VAREdit strikes a much better balance, demonstrating the strength of the VAR paradigm for precise generation. VAREdit also achieves competitive or leading scores on conventional CLIP-based metrics.
Robustness Across Categories
The radar charts in Figure 4 break down performance by editing types. VAREdit achieves state-of-the-art performance across the vast majority of categories. While the 2.2B model shows some limitations in challenging global style and text editing tasks, the 8.4B model substantially closes this performance gap. This illustrates the excellent scaling properties of our framework, suggesting that performance can be further enhanced by scaling to larger models and datasets. Detailed numerical results are presented in the Appendix.
Inference Efficiency
@@
Qualitative Results
Figure 5: Qualitative results among the compared methods and VAREdit of three editing samples in EMU-Edit and PIE-Bench datasets. GPT scores are reported to evaluate adherence to the editing instructions.
Figure 5 provides a visual comparison that reveals the underlying reasons for VAREdit’s quantitative success. In the first example, diffusion-based methods tend to overedit the image and achieve lower GPT-Over. scores. For instance, InstructPix2Pix alters the color of the entire ground and ICEdit erroneously removes the pole. The vanilla AR-based EditAR fails to execute the instruction at all. While this leads to a high GPT-Over. score, it receives a very low GPT-Suc. score. VAREdit successfully completes the task while precisely preserving the unchanged regions, thus achieving the highest GPT-Bal. score. Similar observations hold for the subsequent examples, confirming the effectiveness of VAREdit.
Dataset Setting GPT Score
@@
EMU-Edit Full 5.775 5.892 4.967
Finest 5.819 6.480 5.248
SAR 6.072 7.058 5.565
PIE-Bench Full 7.184 7.170 6.423
Finest 7.291 7.501 6.588
SAR 7.234 7.804 6.684
Table 2: Ablation quantitative results of VAREdit in full-scale, finest-scale and SAR-enhanced settings.
Ablations and Analysis
To isolate the contribution of the SAR module, we conduct an ablation study comparing three conditioning strategies:
  • Full: Conditions on features from all source scales.
  • Finest: Conditions on the finest-scale source features.
  • SAR: Our proposed SAR-enhanced conditions.
The results in Table 2 and Figure 6 confirm our hypothesis on $256\times 256$ resolution 2.2B models. The Full setting yields the lowest GPT-Bal. score, which can be attributed to significantly lower GPT-Over. scores. Introducing all source scales to the conditioning creates distractions for predicting target features, thereby leading to over editing. Moreover, this setting is 60% slower than the other two variants due to longer token sequences. Compared with the Finest setting, the SAR-enhanced model achieves superior GPT-Over. scores, demonstrating the effectiveness of scale-matched information injection. The visual results further support this analysis. The Full and Finest settings introduce unintended textual detail changes and incomplete style references, a flaw that our SAR-based variant avoids.
@@
Conclusion
In this paper, we have proposed VAREdit, an instruction-guided image editing framework following the novel next-scale prediction paradigm in visual autoregressive (VAR) modeling. VAREdit takes the instruction and quantized visual token features into a VAR Transformer model to predict multi-scale residuals of the desired target image, which enjoys enhanced editing adherence and generation efficiency. We have analyzed the efficacy of different conditioning strategies and proposed a novel SAR module to effectively inject scale-matched conditions into the first self-attention layer. Extensive experiments have clearly demonstrated the superior performance of VAREdit, achieving significantly higher editing precision scores and faster generation speed when compared against state-of-the-art methods. As an initial exploration, we hope this study provides valuable new insights into the future design of more effective and efficient AR-based image editing models.

### v2

@@ 3.2.2 Scale Dependency Analysis
The critical scale mismatch challenge introduced by the efficient finest-scale approach raises a fundamental question: which source scales are truly necessary for high-fidelity editing? To investigate such scale dependencies of target residuals and source residuals, we perform a diagnostic analysis of the self-attention mechanisms in a model trained on the full-scale source features. This full-scale setting allows the model to freely attend to all source scales.
@@
Figure 3: Self-attention heatmaps of PIE-Bench samples based on the full-scale transformer across different layers. The dependency patterns differ in the first self-attention layer versus others, inspiring a scale-aligned reference module.
The self-attention heatmaps in Figure 3, averaged across all heads and samples from PIE-Bench (Ju et al., 2023), reveal that different Transformer layers develop distinct attention patterns. In the first self-attention layer, when predicting tokens for a given target scale, the attention mechanism distributes broadly, focusing heavily on the corresponding and all coarser source scales. This pattern indicates that the initial layer is responsible for establishing the global layout and long-range dependencies. This behavior shifts in deeper layers, where attention patterns become highly localized. They exhibit strong diagonal structures, suggesting that attention is primarily confined to tokens in the spatial neighborhood. This functional transition suggests a shift from global structuring to local refinement, where the model copies and adjusts high-frequency details. For this latter task, the fine-grained information provided by $\mathbf{F}_{K}^{(src)}$ is sufficient. This motivates us to design a hybrid solution that provides the scale-aligned reference in the first layer while all subsequent layers attend only to the finest-scale source. See Appendix A.7 for self-attention heatmaps of all model layers.
3.2.3 Scale-Aligned Reference
@@
During generation, when predicting the tokens for target scale $k$, the first self-attention layer computes queries $\mathbf{Q}_{k}^{(tgt)}$ and attends to a combined set of keys and values, which are derived from two sources: (i) the newly generated, scale-aligned reference feature $\mathbf{F}_{k}^{(ref)}$ and (ii) the causal history of previously generated target tokens. Specifically, the self-attention output $\hat{\mathbf{O}}_{k}^{(tgt)}$ is computed as:
@@
$\hat{\mathbf{O}}_{k}^{(tgt)}=\operatorname{Softmax}\left(\frac{\mathbf{Q}_{k}^{(tgt)}\left[\mathbf{K}_{k}^{(ref)\top},\mathbf{K}_{1}^{(tgt)\top},\cdots,\mathbf{K}_{k}^{(tgt)\top}\right]}{\sqrt{d}}\right)\cdot\left[\mathbf{V}_{k}^{(ref)\top},\mathbf{V}_{1}^{(tgt)\top},\cdots,\mathbf{V}_{k}^{(tgt)\top}\right]^{\top},$ (7)
@@
where $(\mathbf{K}_{k}^{(ref)},\mathbf{V}_{k}^{(ref)})$ are projected from the scale-aligned source reference $\mathbf{F}_{k}^{(ref)}$, and $(\mathbf{K}_{1:k-1}^{(tgt)},\mathbf{V}_{1:k-1}^{(tgt)})$ are from previous target scales. This method reduces the spatial size of attention scores before the softmax operation without changing the output sequence length. Crucially, this SAR mechanism is applied only to the first self-attention layer. All subsequent layers operate using only the finest-scale conditioning, attending to $\mathbf{F}_{K}^{(src)}$ and the target’s history residuals. This allows VAREdit to capture crucial scale dependencies in an aligned manner while achieving the efficiency of the finest-scale approach.
4 Experiments
4.1 Experimental Setup
@@
Datasets. VAREdit is trained on a large-scale dataset of 3.92 million paired examples, aggregated from the SEED-Data-Edit (Ge et al., 2024) and ImgEdit (Ye et al., 2025) datasets. From the SEED-Data-Edit dataset, we extracted all single-turn samples and decomposed multi-turn conversations into single-turn editing pairs. These pairs were then filtered using a vision-language model (Team et al., 2025) to remove instances with poor instruction-following quality. Finally, all single-turn samples from ImgEdit were incorporated. Additional details regarding this data processing pipeline are available in Appendix A.1. For evaluation, we assess VAREdit on four established benchmarks: (1) EMU-Edit (Sheynin et al., 2024), including 3,589 samples across eight distinct editing types, (2) PIE-Bench (Ju et al., 2023), including 700 samples that cover ten different editing types, (3) ImgEdit-Bench (Ye et al., 2025), including 737 samples in the basic-edit suite covering nine editing types and 47 complex samples in the understanding-grounding-editing suite, and (4) GEdit-Bench (English subset) (Liu et al., 2025), including 606 samples covering eleven extensive editing types.
@@
Evaluation Metrics. Standard benchmarks like EMU-Edit and PIE-Bench rely on CLIP-based scores (Radford et al., 2021). EMU-Edit employs caption-image similarity (CLIP-Out.) and text-image directional similarity (CLIP-Dir.), whereas PIE-Bench measures similarity for the whole image (CLIP-Whole) and the edited region (CLIP-Edit). However, these metrics often fail to capture important editing quality aspects, such as spurious modifications or incomplete edits. To address these shortcomings, we also adopt the evaluation protocol from OmniEdit (Wei et al., 2025), which uses GPT-4o (Hurst et al., 2024) as an automated judge to provide two key scores on a scale of 0-10: (1) GPT-Success (Suc.): Measures the adherence to the editing instruction (higher is better) and (2) GPT-Overedit (Over.): Assesses the preservation of unedited regions (higher is better). Since a model could achieve a perfect GPT-Over. score by simply disregarding the edit instruction and outputting the original image, we further introduce GPT-Balance (Bal.) as an evaluation metric for overall editing performance, defined as the harmonic mean of the GPT-Suc. and GPT-Over. scores. Detailed prompts and computation methods, along with evaluation details about ImgEdit-Bench and GEdit-Bench, are provided in Appendix A.2.
@@
Compared Approaches. To ensure a comprehensive and rigorous evaluation, we assess VAREdit against several state-of-the-art tuning-based methods. Our comparative analysis includes a broad spectrum of basic open-source methods: InstructPix2Pix (Brooks et al., 2023), UltraEdit (Zhao et al., 2024), OmniGen (Xiao et al., 2025), AnySD (Yu et al., 2025), EditAR (Mu and others, 2025), ACE++ (Mao et al., 2025), ICEdit (Zhang et al., 2025b). We also include four frontier image editing approaches for comparison: GPT-4o-Image (Hurst et al., 2024), Step1X-Edit (Liu et al., 2025), FLUX.1 Kontext (dev) (Labs et al., 2025) and Qwen-Image-Edit (Wu et al., 2025a). The detailed descriptions of these compared approaches are provided in Appendix A.3.
@@
Implementation Details. Our VAREdit models are initialized with weights from the pre-trained Infinity model (Han et al., 2025) along with its bitwise multi-scale residual quantizer (Zhao and others, 2024). To differentiate source image tokens from target image tokens, we introduce a positional offset of $\Delta=(64,64)$ to the 2D Rotary Position Embeddings (2D-RoPE) for all source tokens. To investigate scaling properties, we develop two distinct model sizes: VAREdit-2B and VAREdit-8B. VAREdit-2B undergoes a two-stage training procedure: an initial 8k iterations trained at $256\times 256$ resolution with batch size 1,536 and learning rate 6e56e-5, followed by 7k fine-tuning iterations at $512\times 512$ with batch size 960 and learning rate 1.875e51.875e-5. The larger VAREdit-8B is trained directly at $512\times 512$ resolution for 60k iterations with batch size 1,536 and learning rate 6e56e-5. During training, we optimize the bitwise classifier loss over target residual token indices at each scale, following Infinity (Han et al., 2025). As for inference, we employ a classifier-free guidance (CFG) strength of $\eta=4$ and a logits temperature of $\tau=0.5$. Please refer to Appendix A.4 for more implementation details.
4.2 Quantitative Results
Superior Editing Quality. The quantitative results in Table 1 demonstrate the superiority of VAREdit in editing performance. Among basic open-source methods, VAREdit consistently outperforms all diffusion-based and autoregressive baselines on our primary metric, GPT-Balance. Our VAREdit-8B model achieves a GPT-Bal. score of 7.892 on EMU-Edit and 8.105 on PIE-Bench, surpassing the strongest competitor (ICEdit on EMU-Edit, UltraEdit on PIE-Bench) by 64.9% and 45.3%, respectively. This underscores VAREdit’s ability to perform precise edits while preserving unchanged regions. On conventional CLIP-based metrics, VAREdit also consistently outperforms these baselines and secures the top scores on both EMU-Edit and PIE-Bench. Compared to the four frontier approaches, GPT-4o-Image performs best on both the two benchmarks due to the large scale of training data and model size. Among the open-source approaches, VAREdit-8B performs better than Step1X-Edit and FLUX.1 Kontext (dev) according to GPT-scores and a slightly lower than Qwen-Image-Edit (20B). We have further evaluated the performance of VAREdit on GEdit-Bench. Please refer to Appendix A.6 to access the detailed results and analysis.
Notably, GPT-Bal. serves as a crucial metric since it penalizes a common failure mode: models like OmniGen can achieve high GPT-Over. scores simply by being too conservative and failing to perform the requested edit. To distinguish true preservation from mere inaction, we conduct a more rigorous analysis. We isolate only the most successful edits (GPT-Suc. $\geq$ 9) and re-evaluate GPT-Over. scores on the subsets for all basic approaches. Figure 4 presents the comparisons with two observations. First, VAREdit produces a significantly larger volume of high-quality edits, demonstrating the superior instruction-following capability. Second, and more critically, VAREdit achieves the highest GPT-Over. score within this successful subset, surpassing even OmniGen. This demonstrates VAREdit’s ability to preserve unchanged regions.
Table 1: Quantitative results of VAREdit, leading open-source methods and frontier image editing approaches on EMU-Edit and PIE-Bench benchmarks.
@@ 4.2 Quantitative Results
VAREdit 8.4B 0.282 0.125 8.284 8.461 7.892 0.266 0.233 8.621 8.536 8.105 1.2s
Step1X-Edit 21B 0.280 0.126 7.799 7.843 7.081 0.254 0.226 7.919 8.069 7.351 12.8s
FLUX.1 Kontext 12B 0.284 0.125 7.419 8.800 7.210 0.256 0.225 7.236 8.923 6.998 28.5s
Qwen-Image-Edit 20B 0.286 0.133 9.113 8.598 8.550 0.263 0.232 9.111 8.789 8.567 66.4s
GPT-4o-Image N/A 0.286 0.117 9.517 9.048 9.142 0.266 0.237 9.440 8.866 8.902 N/A
Figure 4: Scatter plots of successful editing samples capacity versus the GPT-Over. scores over the successful editing subsets of basic approaches on EMU-Edit and PIE-Bench benchmarks.
Figure 5: Fine-grained categorical GPT-Suc. and GPT-Bal. scores of VAREdit and basic approaches on EMU-Edit and PIE-Bench benchmarks.
Table 2: Quantitative results of VAREdit and representative approaches on ImgEdit-Bench.
Method Background Extract Attribute Style Remove Replace Add Motion Hybrid UGE
UltraEdit 3.31 2.02 3.01 3.69 1.71 3.13 3.63 3.57 2.33 2.36
Step1X-Edit 3.19 1.87 3.13 4.44 2.61 3.45 3.90 3.43 2.52 3.11
GPT-4o-Image 4.62 2.96 4.26 4.75 3.81 4.49 4.65 4.76 4.54 4.70
VAREdit-2B 2.95 1.94 3.33 4.05 3.51 3.97 3.57 2.88 2.57 2.81
VAREdit-8B 3.93 1.52 3.98 4.60 4.17 4.54 4.06 4.31 3.30 4.09
Robustness Across Categories. We present the radar charts in Figure 5 that break down performance by editing types in EMU-Edit and PIE-Bench. VAREdit achieves state-of-the-art performance across the vast majority of categories among all basic approaches. While VAREdit-2B shows some limitations in challenging global style and text editing tasks, VAREdit-8B substantially closes this performance gap and beats all competitors. This illustrates the excellent scaling properties of our framework, suggesting that performance can be further enhanced by scaling to larger models and datasets. Detailed numerical results are presented in Appendix A.8.
We also present the fine-grained results in ImgEdit-Bench involving all categories in the basic-edit suite and understanding-grounding-editing (UGE) suites. According to the results in Table 2, GPT-4o-Image performs perfectly across almost all categories. Among open-source models, VAREdit-8B performs the best in background change, style transfer, removal, replacement, addition, motion change, hybrid edit and fine-grained UGE. VAREdit achieves satisfactory performance close to or even better than closed-source GPT-4o-Image, demonstrating its superiority across categories.
@@
4.3 Qualitative Results
Figure 6 provides a visual comparison that reveals the underlying reasons for VAREdit’s quantitative success. In the first example, diffusion-based methods tend to overedit the image and achieve lower GPT-Over. scores. For instance, InstructPix2Pix alters the color of the entire ground and ICEdit erroneously removes the pole. The vanilla AR-based EditAR fails to execute the instruction at all, with a high GPT-Over. score yet a very low GPT-Suc. score. VAREdit successfully completes the task while precisely preserving the unchanged regions, thus achieving the highest GPT-Bal. score. Similar observations hold for the subsequent examples, confirming the effectiveness of VAREdit. To access further qualitative results and analysis, please refer to Appendix A.5.
Figure 6: Qualitative results among the compared methods and VAREdit of three editing samples in EMU-Edit and PIE-Bench datasets. GPT scores are also reported to evaluate the editing quality.
4.4 Ablations and Analysis
Table 3: Ablation quantitative results of VAREdit-2B in full-scale, finest-scale and SAR-enhanced settings.
Dataset Setting GPT Score Input Length Time
@@ 4.4 Ablations and Analysis
EMU-Edit Full 5.775 5.892 4.967 1,042 0.8s
Finest 5.819 6.480 5.248 777 0.5s
w/ SAR 6.072 7.058 5.565 777 0.5s
PIE-Bench Full 7.184 7.170 6.423 1,042 0.8s
Finest 7.291 7.501 6.588 777 0.5s
w/ SAR 7.234 7.804 6.684 777 0.5s
To isolate the contribution of the SAR module, we conduct an ablation study comparing the three conditioning strategies for the VAR Transformer: (1) Full: Conditions on features from all source scales, (2) Finest: Conditions on the finest-scale source features, and (3) SAR: Our proposed SAR-enhanced conditions (on the finest-scale source features).
@@
The results in Table 3 and Figure 7 confirm our hypothesis on $256\times 256$ resolution VAREdit-2B. The Full setting yields the lowest GPT-Bal. score, which can be attributed to significantly lower GPT-Over. scores. Introducing all source scales to the conditioning creates distractions for predicting target features and leads to over-editing. Moreover, this setting is 60% slower than the other two variants due to longer token sequences. In contrast, the Finest setting requires fewer tokens as input than the Full setting, yet it suffers from scale mismatch as analyzed in Section 3.2.2 and results in poor performance. Compared with the Finest setting, the SAR-enhanced variant achieves higher GPT-Over. scores, demonstrating the effectiveness of scale-matched information injection. According to the visual results, our SAR-based variant further reduces unintended textual detail changes and incomplete style references, achieving more precise editing results.
5 Conclusion
We have proposed VAREdit, an instruction-guided image editing framework following the next-scale prediction paradigm. VAREdit takes the instruction and quantized visual token features into a VAR Transformer model to predict multi-scale residuals of the desired target image. We have analyzed the efficacy of different conditioning strategies and proposed a novel SAR module to effectively inject scale-matched conditions into the first self-attention layer. Extensive experiments have demonstrated that VAREdit achieves significantly higher editing precision scores and faster generation speed compared to state-of-the-art methods. We hope this initial exploratory study provides valuable insights into the future design of more effective and efficient AR-based visual models.
Acknowledgement. This work was supported by the Key Science & Technology Project of Anhui Province No. 202523o09050002, the National Natural Science Foundation of China (No. 62337001), the Key Technologies R & D Program of Anhui Province (No. 202423k09020039) and the Fundamental Research Funds for the Central Universities.